Solve to find all the roots of the following equation:
4w^2 = -2w + 6

Answer by checkley71(8403)

4W~2+2W-6=0 OR 2W~2+X-3=0 OR (2W+3)(W-1)=0 OR 2W+3=0 OT 2W=-3 OR W=-3/2
W-1=0 OR W=1